OpenStudy (anonymous):

is a negetive 3/7 equal to a -3/-7 equal to 3/7

OpenStudy (anonymous):

no , \[-\left( \frac{ p }{ q } \right)=\frac{ -p }{ q }\]

OpenStudy (anonymous):

or \[-\left( \frac{ p }{ q } \right)=\frac{ p }{ -q }\]

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

you're welcome.

OpenStudy (anonymous):

wonderwoman should know this!! che che che

OpenStudy (anonymous):

no.......-3/-7=3/7 but not -3/7

OpenStudy (anonymous):

well wonderwoman don't

OpenStudy (anonymous):

\[\frac{ -3 }{ -7 }=\frac{ -1*3 }{ -1*7 }=\frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

shame on u amazon!